# Multi: Multimodal Understanding Leaderboard with Text and Images

### v3

@@ Abstract
Multi: Multimodal Understanding Leaderboard with Text and Images
Zichen Zhu1and Yang Xu1and Lu Chen [^2] The corresponding authors are Lu Chen and Kai Yu.12and Jingkai Yang1and Yichuan Ma1 Yiming Sun1and Hailin Wen1and Jiaqi Liu1and Jinyu Cai1 Yingzi Ma1and Situo Zhang1and Zihan Zhao1and Liangtai Sun1and Kai Yu [^2] The corresponding authors are Lu Chen and Kai Yu.12 1 X-LANCE Lab, Department of Computer Science and Engineering MoE Key Lab of Artificial Intelligence, SJTU AI Institute Shanghai Jiao Tong University, Shanghai, China 2 Suzhou Laboratory, Suzhou, China {JamesZhutheThird, xuyang0112, chenlusz, kai.yu}@sjtu.edu.cn
@@ 1 Introduction
The rapid development of large-scale language models (LLMs) Sun et al. (2023); Chen et al. (2022); qwe (2024); Cai et al. (2024); OpenAI (2022, 2023a); Team (2023); Touvron et al. (2023) has brought about remarkable progress in natural language processing (NLP) and related fields. However, human communication and understanding extend far beyond just language, encompassing various other forms of information such as images, mathematical formulas, graphs, tables, and diagrams. These visual modalities play an essential role in fields like science and engineering, where they are crucial for conveying complex ideas. Consequently, there has been increasing interest in developing Multimodal Large Language Models (MLLMs) Du et al. (2022); Hu et al. (2023); LinkSoul-AI (2023); Bai et al. (2023a); 01.ai (2023); Chen et al. (2024); Zhe et al. (2024); Yao et al. (2024); Wang et al. (2024); OpenAI (2023b); Team (2023); OpenAI (2024a, b); Anthropic , that can process and generate across different modalities, including visual ones, and perform tasks that require cross-modal reasoning. All these rapid advancements raise the question of how a model compares to a real student.
@@
Evaluating MLLMs presents unique challenges, as current benchmarks Lu et al. (2022); Li et al. (2024a) are either too narrow in scope, focusing on natural scene images, or too simplistic to effectively assess the capabilities of modern models. Many existing scientific benchmarks Sun et al. (2024); Huang et al. (2024) rely on multiple-choice questions with a single correct answer, which may not adequately test a model’s comprehension or reasoning skills. As a result, models can sometimes arrive at superficial solutions without fully understanding the problem. A more robust, detailed, and multi-scale dataset is necessary to evaluate MLLMs in a more comprehensive manner under diverse real-world conditions. Additionally, while most existing benchmarks are in English, the rapid progress of Chinese multimodal models has highlighted the need for a Chinese multimodal benchmark that includes both Chinese text and visual content, offering unique challenges to the field.
@@
In this paper, we introduce Multi, a new multimodal benchmark for evaluating LLMs on cross-modal understanding tasks. Our primary goal was to evaluate MLLMs across a broad range of tasks that a typical Chinese student would encounter throughout their academic progression. The Multi benchmark, which stands for Multimodal Understanding Leaderboard with Text and Images, consists of 18,430 carefully curated questions derived from real-world educational and online sources. These questions span multiple scientific domains, including mathematics, physics, and computer science, and feature a variety of question formats, such as multiple-choice, fill-in-the-blank, and open-ended questions. Importantly, Multi also includes driving and administrative aptitude tests, which are significant in the Chinese context and offer further challenges in cross-modal reasoning and generation.
@@
To push the boundaries of MLLM evaluation, we also introduce two specialized subsets within Multi: Multi-Elite, which includes 500 tough questions designed to test the limits of current MLLMs, and Multi-Extend, a set of 4,596 external knowledge pieces that assess models’ abilities in in-context learning and knowledge transfer. These subsets provide deeper insights into the strengths and weaknesses of MLLMs, driving further research and model improvement. Sample questions from Multi are presented in Figure 2, and additional details are available in E and F.
@@ 2 Related Works
With advancements in aligning features across multiple modalities, like CLIP Radford et al. (2021) and ALBEF Li et al. (2021), recent studies have explored projecting vision features into the latent space of LLMs, aiming to enhance their capabilities of comprehending visual information. For example, BLIP-2 Li et al. (2023a) pioneers this approach by employing Q-Former to translate image features into text representations. Following this, LLaVA Liu et al. (2024), MiniGPT-4 Zhu et al. (2024), and InstructBLIP Dai et al. (2024) have introduced visual instruction tuning to bolster the capability of MLLMs of following instructions. Our primary focus is on the proficiency of MLLMs in comprehending instructions in Chinese, which are divided into two main branches: open-source models, which typically build upon existing Chinese LLMs or are fine-tuned on Chinese instruction datasets, examples of which include Chinese-LLaVA LinkSoul-AI (2023), VisualGLM Du et al. (2022), VisCPM Hu et al. (2023), Qwen-VL Bai et al. (2023a), InternVL Zhang et al. (2023), Yi-VL 01.ai (2023); and closed-source models, which are often highly powerful, multi-lingual systems such as GPT-4V(ision) OpenAI (2023b) and Gemini Team (2023). In this paper, we intend to evaluate these models across a range of scientific fields on the Multi benchmark, offering an extensive assessment and guidance for the onward trajectory of Chinese MLLMs.
@@
In assessing MLLMs, traditional methods primarily rely on established vision-language (VL) benchmark datasets. Renowned benchmarks such as VQA Goyal et al. (2017), OK-VQA Antol et al. (2015), GQA Hudson and Manning (2019), and MSCOCO Lin et al. (2014) are tailored to specific VL tasks like image captioning, open-domain visual question answering, and visual reasoning. While the evaluation based on standard benchmark datasets yields significant insights into MLLMs’ capabilities, these approaches may not entirely capture their comprehensive intelligence in real-world scenarios. Therefore, a diverse array of benchmarks has been developed to examine MLLMs in dealing with various tasks in the real world. Benchmarks like LLaVA-Bench Liu et al. (2024), MMBench Liu et al. (2025), MM-VET Yu et al. (2024), TouchStone Bai et al. (2023b), MLLM-bench Ge et al. (2023), and SEED-BenchLi et al. (2024a, b), for instance, leverage GPT to evaluate the relevance and helpfulness of human-like long responses in the reality. POPE Li et al. (2023b) and HallusionBench Guan et al. (2024) introduce various analytical criteria for the holistic evaluation of MLLMs’ hallucinations. Furthermore, M3Exam Zhang et al. (2024b), SciGraphQA Li and Tajbakhsh (2023), MathVista Lu et al. (2024), AGIEval Zhong et al. (2024), and MMMU Yue et al. (2024) consider MLLMs as experts to extend the evaluation scope by incorporating advanced perception and reasoning within domain-specific knowledge, for example, scientific questions and driving tests. The works most related to us are M3Exam, ScienceQA, SciEval Sun et al. (2024) and C-Eval Huang et al. (2024). Our approach distinguishes itself by offering a broader spectrum of question types compared to the first two and supports a multimodal evaluation in contrast to the last two.
@@ 3 The Multi Benchmark
We propose Multi, a Multimodal Understanding Leaderboard with Text and Images, which can serve as a challenging and diverse benchmark for the MLLM community. In § 3.1, we provide an overall introduction to the data construction process, with more details provided in E. The detailed statistics are provided in § 3.2. In § 3.3 and § 3.4 we introduce Multi-Elite and Multi-Extend. We also give a comparison between Multi and other knowledge-based benchmarks in § 3.5
@@ 3.1 Data Construction Process
Figure 3: The construction pipeline of Multi
The data construction pipeline is shown in Figure 3. To develop Multi, we follow several key steps to ensure high-quality and precise annotation. Firstly, we crawl open-source raw question data from the Internet and transcript close-source exams from paper documents. Secondly, we format each question and knowledge piece into markdown and  formula format to maintain precision and quality. Thirdly, we revise and refine each question multiple times to prevent data leakage and increase difficulty. Lastly, We rate every question based on its difficulty and content richness.
@@
We collect more than 2.7M raw data from the Internet, ranging from exams and quizzes from Chinese junior and senior schools and several society exams. We design an algorithm to pick out a proportion of the questions as the fundamental data of our benchmark. The selection is based on the questions’ text length, number of images, corresponding subjects, and knowledge pieces, to reach a higher diversity of questions and coverage of knowledge. The details are presented in D. We also collect questions from internal exams and practices of several top universities. After the selection, we obtain over 18K questions as the raw data.
@@
The data process and annotation for our dataset involve a comprehensive series of steps to ensure high-quality, diverse content.
@@
Throughout these stages, we process 2.7 million questions in total and pick out 18,430, incorporating 23,320 scoring points, 7,658 images, and 4,595 knowledge pieces. Multi highlights a broad diversity in question types, including multiple-choice questions with both single and multiple answers, along with fill-in-the-blank and open-ended writing questions enriching the testing scenarios.  [^1] For the sake of simplifying writing, in the following paragraphs we may use abbreviations. We denote multiple-choice questions with a single answer as SA or Single Answer Choosing and those with multiple answers as MA or Multi Answer Choosing. We use FB for fill-in-the-blank questions and OP for open-ended writing questions. The stages during data processing and annotation significantly increase the diversity and difficulty of the dataset. For details of data processing and annotation, please refer to E.
@@ 3.2 Statistics
We provide detailed statistics in Table 1. We also provide the composition of education level and subjects as shown in Figure 4.
@@
Open Ended Writing 797(4.33%) 797(3.48%)
@@
Figure 4: The question distribution by education level and subjects
@@
Table 2: The answer distribution for multiple-choice questions
@@ 3.3 The Multi-Elite Set
We have selected 500 objective questions to create a hard subset. These multiple-choice and fill-in-the-blank questions are more evenly distributed across various educational levels and subjects and have been evaluated by annotators as high in difficulty and quality, featuring detailed content and complex reasoning requirements. In the selection process, the performance of several early-released models during earlier evaluation is also taken into consideration. We especially focused on poorly answered questions by GPT-4VOpenAI (2023b), the state-of-the-art model at the time. As a result, these models may get lower scores compared to the random baseline in Table 7. Since a complete run on Multi could be costly for some models, we encourage researchers to run their experiments on these 500 high-quality questions.
@@ 3.4 The Multi-Extend Background Knowledge Dataset
External knowledge is crucial to provide critical information that assists in solving questions using the In-Context Learning (ICL) abilities. Some of the raw questions retrieved from the Internet have corresponding knowledge pieces attached. We also collect more knowledge pieces for uncovered questions with the assistance of LLMs and outer knowledge sources (e.g. New Bing [^2] https://bing.com/new and Wikipedia [^3] https://wikipedia.org). We conduct annotations on these knowledge pieces to confirm the correctness of the content and present them in the Multi-Extend dataset. This dataset consists of about 4.6K knowledge pieces, designed to test the in-context learning abilities and knowledge transfer skills of models. This dataset provides comprehensive insights into the capabilities and limitations of multimodal LLMs, opening new pathways for research exploration.
@@ 3.5 Comparison with Existing Benchmarks
VQA Antol et al. (2015) en 36 764K - 265K - Repurposed
ScienceQA Lu et al. (2022) en 21 21K 19K 10K 0.3K Textbooks
SciBench Wang et al. (2023) en 6 0.8K - 0.1K - Textbooks
M3Exam Zhang et al. (2024b) 9 langs 4 12K - 3.1K - Exams
AGIEval Zhong et al. (2024) zh, en 20 8K a few - - Exams
CMMU He et al. (2024) zh 7 4K - 4K - Exams
GAOKAO-MM Zong and Qiu (2024) zh 8 0.6K - 0.9K - Exams
MMBench Liu et al. (2025) zh/en 20 3K - 3K - Web, Repurposed
SEED-Bench Li et al. (2024a) en 12 19K - 19K+ - Anno.
SEED-Bench-2 Li et al. (2024b) en 27 24K - 22K+ - Anno.
MLLM-Bench Ge et al. (2023) en 42 0.4K - 0.4K - Anno.
Touchstone Bai et al. (2023b) en 27 0.9K - 0.9K - Anno.
AlignMMBench Wu et al. (2024) zh 13 5K - 1K - Web, Generated
C-Eval Huang et al. (2024) zh 52 14K a few - - Exams, Web
SciEval Sun et al. (2024) en 3 18K - - - Web, Repurposed
MMMU Yue et al. (2024) en 30 12K 2K 11K+ - Anno., Web, Textbooks
CMMMU Zhang et al. (2024a) zh 30 12K - 12K - Anno., Web, Textbooks
@@
We list the features of existing benchmarks and make a comparison with Multi in Table 3. We believe that Multi assembles the most advantages of the existing benchmarks and is sure to provide a good option for the community to test the capabilities of their Vision LLMs.
@@ 4 Experiments
In this section, we conduct a comprehensive experiment and several ablation studies. We firstly introduce baselines, settings, and metrics in § 4.1, § 4.2 and § 4.3. The main experiment analysis is in § 4.4, followed by several research questions and ablation study results from § 4.5 to § 4.10. We conclude our findings and thoughts as take-home messages in § 4.11
@@ 4.1 Baselines
We evaluate a wide range of leading-edge MLLMs that support Chinese, including both open-sourced models LinkSoul-AI (2023); Bai et al. (2023a); Wang et al. (2024); Hu et al. (2023); Du et al. (2022); Chen et al. (2024); 01.ai (2023) and closed-source models OpenAI (2024a, 2023b); Team (2023); Anthropic ; OpenAI (2024b). We evaluate these models with both multimodal input and text-only input to verify the information gained from input images. We also select several most capable and representative LLMs Chen et al. (2022); Sun et al. (2023); OpenAI (2022, 2023a); qwe (2024); Zhang et al. (2023); Cai et al. (2024) for comparison. We also add ablation studies to verify the improvement by giving the OCR information or captions to these LLMs. For detailed information about all models evaluated in Multi, please refer to Table 10 in B.
@@
Seven human experts from top universities in China are involved in the human evaluation. The complete human evaluation set includes all 500 questions from Multi-Elite and another 500 randomly sampled questions from Multi. Each question was assigned to two experts based on the subjects they had chosen in Gaokao. We follow the rule in MMMUYue et al. (2024) where the experts are allowed to consult their textbooks and Multi-Extend but are forbidden to search the Internet. We calculate the score of Multi as the weighted average of each subject.
@@
We also provide a random baseline and a most frequent baseline. For the random baseline, we randomly choose one choice in SA, randomly select/drop each choice in MA, and randomly choose from [A, B, C, D, 1, 2, 3, 4, True, False] in OP, and simply take the question content as the answer.
Figure 5: An example of the prompts used when evaluating a multiple-choice question with image context, knowledge piece, and single correct answer.

### v4

@@ Abstract
keywords:
Multimodal, Large Language Model, Logic Reasoning, Image Comprehension, Benchmark
@@ 1 Introduction
The rapid development of large-scale language models (LLMs) [1, 2, 3, 4, 5, 6, 7, 8] has brought about remarkable progress in natural language processing (NLP) and related fields [9, 10, 11, 12, 13, 14]. However, human communication and understanding extend far beyond just language, encompassing various other forms of information such as images, mathematical formulas, graphs, tables, and diagrams. These visual modalities play an essential role in fields like science and engineering, where they are crucial for conveying complex ideas. Consequently, there has been increasing interest in developing Multimodal Large Language Models (MLLMs) [15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 7, 25, 26, 27], that can process and generate across different modalities, including visual ones, and perform tasks that require cross-modal reasoning [28, 29, 30]. All these rapid advancements raise the question of how a model compares to a real student.
@@
Evaluating MLLMs presents unique challenges, as current benchmarks [31, 32] are either too narrow in scope, focusing on natural scene images, or too simplistic to effectively assess the capabilities of modern models. Many existing scientific benchmarks [33, 34] rely on multiple-choice questions with a single correct answer, which may not adequately test a model’s comprehension or reasoning skills. As a result, models can sometimes arrive at superficial solutions without fully understanding the problem. A more robust, detailed, and multi-scale dataset is necessary to evaluate MLLMs in a more comprehensive manner under diverse real-world conditions. Additionally, while most existing benchmarks are in English, the rapid progress of Chinese multimodal models has highlighted the need for a Chinese multimodal benchmark that includes both Chinese text and visual content, offering unique challenges to the field.
@@
In this paper, we introduce Multi, a new multimodal benchmark for evaluating LLMs on cross-modal understanding tasks. Our primary goal is to evaluate MLLMs across a broad range of tasks that a typical Chinese student would encounter throughout their academic progression, as shown in Figure 1. The Multi benchmark, which stands for Multimodal Understanding Leaderboard with Text and Images, consists of 18,430 carefully curated questions derived from real-world educational and online sources. These questions span multiple scientific domains, including mathematics, physics, and computer science, and feature a variety of question formats, such as multiple-choice, fill-in-the-blank, and open-ended questions. Importantly, Multi also includes driving and administrative aptitude tests, which are significant in the Chinese context and offer further challenges in cross-modal reasoning and generation.
@@
To push the boundaries of MLLM evaluation, we also introduce two specialized subsets within Multi: Multi-Elite, which includes 500 tough questions designed to test the limits of current MLLMs, and Multi-Extend, a set of 4,596 external knowledge pieces that assess models’ abilities in in-context learning and knowledge transfer. These subsets provide deeper insights into the strengths and weaknesses of MLLMs, driving further research and model improvement. Sample questions from Multi are presented in Figure 2, and additional details are available in F and G.
@@ 2 Related Work
With advancements in aligning features across multiple modalities, like CLIP [35] and ALBEF [36], recent studies have explored projecting vision features into the latent space of LLMs, aiming to enhance their capabilities of comprehending visual information. For example, BLIP-2 [37] pioneers this approach by employing Q-Former to translate image features into text representations. Following this, LLaVA [38], MiniGPT-4 [39], and InstructBLIP [40] have introduced visual instruction tuning to bolster the capability of MLLMs of following instructions. Our primary focus is on the proficiency of MLLMs in comprehending instructions in Chinese, which are divided into two main branches: open-source models, which typically build upon existing Chinese LLMs or are fine-tuned on Chinese instruction datasets, examples of which include Chinese-LLaVA [17], VisualGLM [15], VisCPM [16], Qwen-VL [18], InternVL [41], Yi-VL [19]; and closed-source models, which are often highly powerful, multi-lingual systems such as GPT-4V(ision) [24] and Gemini [7]. In this paper, we intend to evaluate these models across a range of scientific fields on the Multi benchmark, offering an extensive assessment and guidance for the onward trajectory of Chinese MLLMs.
@@
In assessing MLLMs, traditional methods primarily rely on established vision-language (VL) benchmark datasets. Renowned benchmarks such as VQA [42, 43], OK-VQA [44], GQA [45], and MSCOCO [46] are tailored to specific VL tasks like image captioning, open-domain visual question answering, and visual reasoning. While the evaluation based on standard benchmark datasets yields significant insights into MLLMs’ capabilities, these approaches may not entirely capture their comprehensive intelligence in real-world scenarios. Therefore, a diverse array of benchmarks has been developed to examine MLLMs in dealing with various tasks in the real world. Benchmarks like LLaVA-Bench [38], MMBench [47], MM-VET [48], TouchStone [49], MLLM-bench [50], and SEED-Bench[32, 51], for instance, leverage GPT to evaluate the relevance and helpfulness of human-like long responses in the reality. POPE [52] and HallusionBench [53] introduce various analytical criteria for the holistic evaluation of MLLMs’ hallucinations. Furthermore, M3Exam [54], SciGraphQA [55], MathVista [56], AGIEval [57], and MMMU [58] consider MLLMs as experts to extend the evaluation scope by incorporating advanced perception and reasoning within domain-specific knowledge, for example, scientific questions and driving tests. The works most related to us are M3Exam, ScienceQA, SciEval [33], and C-Eval [34]. Our approach distinguishes itself by offering a broader spectrum of question types compared to the first two and supports a multimodal evaluation in contrast to the last two.
@@ 3 The M ULTI  Benchmark
We propose Multi, a Multimodal Understanding Leaderboard with Text and Images, which can serve as a challenging and diverse benchmark for the MLLM community. In § 3.1, we provide an overall introduction to the data construction process, with more details provided in F. The detailed statistics are provided in § 3.2. In § 3.3 and § 3.4 we introduce Multi-Elite and Multi-Extend. We also give a comparison between Multi and other knowledge-based benchmarks in § 3.5.
@@ 3.1 Data Construction Process
The data construction pipeline is shown in Figure 3. To develop Multi, we follow several key steps to ensure high-quality and precise annotation. Firstly, we crawl open-source raw question data from the Internet and transcribe closed-source exams from paper documents. Secondly, we format each question and knowledge piece into markdown and  formula format to maintain precision and quality. Thirdly, we revise and refine each question multiple times to prevent data leakage and increase difficulty. Lastly, we rate every question based on its difficulty and content richness.
Figure 3: The construction pipeline of M ULTI . Upper half: The construction of Multi and Multi-Elite. (1) Collect large-scale open-source data. (2) Select high-quality and evenly distributed open-source Q&As. (3) Collect and transcribe questions from closed-source exams. (4) Convert raw data into markdown format using scripts. (5) Skilled annotators refine, split, and label the data. (6) Assess the difficulty and quality of each question. (7) Further refine and increase difficulty through expert review. (8) Select the most challenging items for the Multi-Elite subset. Lower half: The construction of Multi-Extend. (9) Summarize knowledge from open-source data. (10) Generate additional knowledge via LLMs.
@@
We collect more than 2.7M raw data from the Internet, ranging from exams and quizzes from Chinese junior and senior schools and several society exams. We design an algorithm to pick out a proportion of the questions as the fundamental data of our benchmark. The selection is based on the questions’ text length, number of images, corresponding subjects, and knowledge pieces, to reach a higher diversity of questions and coverage of knowledge. The details are presented in E. We also collect questions from internal exams and practices of several top universities. After the selection, we obtain over 18K questions as the raw data.
@@
The data processing and annotation for our dataset involve a comprehensive series of steps to ensure high-quality, diverse content.
@@
Throughout these stages, we process 2.7 million questions in total and pick out 18,430, incorporating 23,320 scoring points, 7,658 images, and 4,595 knowledge pieces. Multi highlights a broad diversity in question types, including multiple-choice questions with both single and multiple answers, along with fill-in-the-blank and open-ended writing questions, enriching the testing scenarios.  [^1] For the sake of simplifying writing, in the following paragraphs we may use abbreviations. We denote multiple-choice questions with a single answer as SA or Single Answer Choosing and those with multiple answers as MA or Multi Answer Choosing. We use FB for fill-in-the-blank questions and OP for open-ended writing questions. The stages during data processing and annotation significantly increase the diversity and difficulty of the dataset. For details of data processing and annotation, please refer to F.
@@ 3.2 Statistics
We provide detailed statistics in Table 2. We also provide the composition of education level and subjects as shown in Table 2.
@@
Open-ended Writing 797(4.33%) 797(3.48%)
@@
Figure 4: The question distribution by the education level and subjects.
@@
Table 2: The answer distribution for multiple-choice questions.
@@ 3.3 The M ULTI -E LITE  Set
We have selected 500 objective questions to create a hard subset. These multiple-choice and fill-in-the-blank questions are more evenly distributed across various educational levels and subjects and have been evaluated by annotators as high in difficulty and quality, featuring detailed content and complex reasoning requirements. In the selection process, the performance of several early-released models during earlier evaluation is also taken into consideration. We especially focused on poorly answered questions by GPT-4V[24], the state-of-the-art model at the time. As a result, these models may get lower scores compared to the random baseline in Table 7. Since a complete run on Multi could be costly for some models, we encourage researchers to run their experiments on these 500 high-quality questions.
@@ 3.4 The M ULTI -E XTEND  Background Knowledge Dataset
External knowledge is crucial to provide critical information that assists in solving questions using the In-Context Learning (ICL) abilities. Some of the raw questions retrieved from the Internet have corresponding knowledge pieces attached. We also collect more knowledge pieces for uncovered questions with the assistance of LLMs and external knowledge sources (e.g., New Bing [^2] https://bing.com/new and Wikipedia [^3] https://wikipedia.org). We conduct annotations on these knowledge pieces to confirm the correctness of the content and present them in the Multi-Extend dataset. This dataset consists of about 4.6K knowledge pieces, designed to test the in-context learning abilities and knowledge transfer skills of models. This dataset provides comprehensive insights into the capabilities and limitations of multimodal LLMs, opening new pathways for research exploration.
@@ 3.5 Comparison with Existing Benchmarks
VQA [42] en 36 764k - 265k - Repurposed
ScienceQA [31] en 21 21k 19k 10k 0.3k Textbooks
SciBench [59] en 6 0.8k - 0.1k - Textbooks
M3Exam [54] 9 langs 4 12k - 3.1k - Exams
AGIEval [57] zh, en 20 8k a few - - Exams
CMMU [60] zh 7 4k - 4k - Exams
GAOKAO-MM [61] zh 8 0.6k - 0.9k - Exams
MMBench [47] zh/en 20 3k - 3k - Web, repurposed
SEED-Bench [32] en 12 19k - 19k+ - Anno.
SEED-Bench-2 [51] en 27 24k - 22k+ - Anno.
MLLM-Bench [50] en 42 0.4k - 0.4k - Anno.
Touchstone [49] en 27 0.9k - 0.9k - Anno.
AlignMMBench [62] zh 13 5k - 1k - Web, generated
C-Eval [34] zh 52 14k a few - - Exams, web
SciEval [33] en 3 18k - - - Web, repurposed
MMMU [58] en 30 12k 2k 11k+ - Anno., web, textbooks
CMMMU [63] zh 30 12k - 12k - Anno., web, textbooks
@@
We list the features of existing benchmarks and make a comparison with Multi in Table 3. We believe that Multi assembles the most advantages of the existing benchmarks and is sure to provide a good option for the community to test the capabilities of their MLLMs.
@@ 4 Experiments
In this section, we conduct a comprehensive experiment and several ablation studies. We firstly introduce baselines, settings, and metrics in § 4.1, § 4.2, and § 4.3. The main experiment analysis is in § 4.4, followed by several research questions and ablation study results from § 4.5 to § 4.10. We conclude our findings and thoughts as take-home messages in § 4.11.
@@ 4.1 Baselines
We evaluate a wide range of leading-edge MLLMs that support Chinese, including both open-sourced models [17, 18, 23, 16, 15, 20, 19] and closed-source models [25, 24, 7, 27, 26]. We evaluate these models with both multimodal input and text-only input to verify the information gained from input images. We also select several most capable and representative LLMs [2, 1, 5, 6, 3, 41, 4] for comparison. We also add ablation studies to verify the improvement by giving the OCR information or captions to these LLMs. For detailed information about all models evaluated in Multi, please refer to Table 10 in C.
@@
Seven human experts from top universities in China are involved in the human evaluation. The complete human evaluation set includes all 500 questions from Multi-Elite and another 500 randomly sampled questions from Multi. Each question was assigned to two experts based on the subjects they had chosen in the Gaokao. We follow the rule in MMMU[58] where the experts are allowed to consult their textbooks and Multi-Extend but are forbidden to search the Internet. We calculate the score of Multi as the weighted average of each subject.
@@
We also provide a random baseline and a most-frequent baseline. For the random baseline, we randomly choose one choice in SA, randomly select/drop each choice in MA, and randomly choose from [A, B, C, D, 1, 2, 3, 4, True, False] in OP, and simply take the question content as the answer.
Figure 5: An example of the prompt template. It is used when evaluating a multiple-choice question with image context, knowledge piece, and single correct answer.